(a) Find the value of m if the angle between a(3,m) and b (2,-1)is 360 degrees

/ 3marks


(b) Find the longest side of right triangle whose other sides is:(√2+1)cm and (√2-1)cm long



1
Expert's answer
2021-09-06T13:31:16-0400

Solution:

(a)

Given:  a(3,m) and b (2,-1)

Using dot product formula of vectors:

"a.b=|a||b|\cos \theta"

"\\\Rightarrow 3(2)+m(-1)=\sqrt{3^2+m^2}\times\sqrt{2^2+(-1)^2}\times \cos(360\degree)
\\\Rightarrow 6-m=\sqrt{(9+m^2)5}\times(1)
\\\Rightarrow (6-m)^2=5(9+m^2)
\\\Rightarrow 36+m^2-12m=45+5m^2
\\\Rightarrow 4m^2+12m+9=0
\\\Rightarrow 4m^2+6m+6m+9=0
\\\Rightarrow (4m+3)(2m+3)=0
\\\Rightarrow m=-\dfrac34,-\dfrac32"

(b)

Two sides are :p=(√2+1)cm and b=(√2-1) cm

We need to find the Hypotenuse (h).

Using Pythagoras theorem,

"h^2=p^2+b^2"

"\Rightarrow h^2=(\sqrt2+1)^2+(\sqrt2-1)^2
\\\Rightarrow h^2=2+1+2\sqrt2+2+1-2\sqrt2
\\\Rightarrow h^2=6
\\\Rightarrow h=\sqrt6\ cm"